figures/purdue Phase 1: Constructing a map of aging pathways in yeast Phase 2: Projecting yeast aging pathways to human tissues

Summary

A Systems Study of Aging and Age-Related Pathologies

Ananth Grama

Center for Science of Information Purdue University

BioKDD, 2013

Grama et al.  Computational Aging